A wooden boat is found buried beneath the ocean floor. Scientists believe that timbers making up the boat had 16 g of carbon-14 in them when alive. They now have 4.1 g of carbon-14. How old is the boat if the half-life of carbon-14 is 5715 years?
1
Expert's answer
2020-05-05T18:40:34-0400

Determine the decimal amount remaining: "\frac{4.1}{16}=0.25625".

Determine the number of half-lives: "(0.5)^n=0.25625",

"n \ln 0.5=\ln 0.25625,\,\,n=\frac{\ln 0.25625}{\ln 0.5}=1.964376".

Determine years elapsed: "5715 \cdot n=11226.4" years.